if the effective mass of electron is 0.4mo having momentum 2.416*10-16 gram cm/sec then calculate energy of free electron in electron voult at the bottom of conduction brand?
1
Expert's answer
2020-06-25T09:49:00-0400
"K=\frac{p^2}{2m}=\frac{(2.416\cdot10^{-16})^2}{2(0.5)(9.109\cdot10^{-31})}=0.064\ J"